OpenStudy (anonymous):

simplifying f(x+h)−f(x)/h where 2x+3

OpenStudy (anonymous):

If it helps you at all, the answer turns out to be 2

OpenStudy (anonymous):

since it is the same as asking for the slope of the line. you are asked for the change in \(y\) over the change in \(x\) which is the same as asking for the slope the slope of your line is 2

OpenStudy (anonymous):

i'm not asking for the slope

OpenStudy (anonymous):

yes, just in disguise

OpenStudy (anonymous):

f(x)= 2x+3

OpenStudy (anonymous):

right and \(f(x+h)-f(x)\) is the change in \(y\)

OpenStudy (anonymous):

no... it says to simplify. straight forward simplifying where you find f(x) and they give you that f(x) = 2x+3

OpenStudy (anonymous):

no no... it is not a slope problem.

OpenStudy (anonymous):

\[f(x+h)-f(x)=2(x+h)+3-(2x+3)=2x+2h+3-2x-3=2h\]

OpenStudy (anonymous):

divide by \(h\) and you get \(2\)

OpenStudy (anonymous):

I like how he's arguing with you that you're not finding the right answer, he hasn't learned that's what it really means yet! :P

OpenStudy (anonymous):

just like computing \(\frac{y_2-y_1}{x_2-x_1}\) with \(y_2=f(x+h), y_1=f(x), (x_2-x_1)=h\)

OpenStudy (anonymous):

no i have the paper with me... and i'm a she. it doesn't say slope anywhere? it says to simplify.